PHOSPHORUS HYPERFINE STRUCTURE IN THE ELECTRONIC SPECTRUM OF THE HPCl FREE RADICAL

\underline{BRANDON S. TACKETT} and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055; ALAN G. ADAM and SCOTT A. SHEPARD, Department of Chemistry, University of New Brunswick, Fredericton, NB., Canada E3B 6E2.

The 444 nm 2_{0}^{1} bands of the Ã^{2}A′ - X̃^{2}A′′ transition of the jet-cooled HP^{35}Cl and HP^{37}Cl radicals have been studied at high resolution using the pulsed electric discharge technique with a precursor mixture of PCl_{3} and H_{2}. Spectra recorded with linewidths of 360 MHz revealed resolved hyperfine structure in both isotopomers arising from the excited state Fermi contact interaction of the unpaired electron with the magnetic moment of the ^{31}P nucleus, with a′_{F} = 0.0641(10) cm^{−1} and 0.0636(31) cm^{−1} for HP^{35}Cl and HP^{37}Cl, respectively. No contribution from the ground state, or excited state contributions from the hydrogen or chlorine nuclei were resolved, confirming ab initio predictions that HPCl is a pπ radical in the X̃ state, and an sσ radical with a substantial contribution from the phosphorus 3s atomic orbital in the Ã state. The free atom comparison method has been used to estimate that the singly occupied molecular orbital in the excited state has 14% phosphorus 3s character.